# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v1

@@ Appendix B Implementation Details
We follow Malladi et al. (2023) and provide the memory-efficient version of DP-ZO in Algorithm 2. Algorithm 2 enjoys the benefit that it does not incur additional GPU memory cost compared to inference.
Algorithm 2 Differentially Private-ZO (GPU memory efficient version. Adapted from Malladi et al. (2023))
@@
2:procedure DP-ZO($\theta$, $\mathcal{D}$, $\varepsilon$, $\sigma$, $T$, $s$, $\phi$, $C$, $\alpha$)
@@
17:procedure PerturbParameters($\theta,\phi,s$)
18:     Reset random number generator with seed $s$
19:     for $\theta_{i}\in\theta$ do
20:         $z\sim\mathcal{N}(0,1)$
21:         $\theta_{i}\leftarrow\theta_{i}+\phi z$
@@ Appendix C Design Choices
$g_{i}$ $=L(\theta+\varepsilon z_{i};B)-L(\theta-\varepsilon z_{i};B)\text{(projected gradient for each }i)$
@@
Algorithm 1 estimates the gradients once. As outlined above, SPSA can be extended to n-SPSA gradient estimator and n-SPSA can improve the performance in the non-private setting (Malladi et al., 2023). Here we discuss our design choice of why we choose $n=1$ in Algorithm 1.
  • Estimate the average. Previous work (Malladi et al., 2023) shows that averaged estimation helps the non-private setting. In a private setting, we have to privatize the gradient estimation. Here we discuss our initial design of the privatized n-SPSA gradient estimation. For the sampled batch, assuming we are adding the Gaussian noise $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ for 1-SPSA. Then for n-SPSA, to ensure we have the same privacy cost as 1-SPSA, we need to add $\mathcal{N}(0,n\cdot\mathrm{C}^{2}\sigma^{2})$ to each gradient estimation and finally average the $n$ gradients. Our privacy analysis follows the $n$-fold composition of Gaussian mechanism (Corollary 3.3 in Gaussian differential privacy (Dong et al., 2019)). Our initial experiment result shows that our current analysis for n-SPSA noise addition does not make n-SPSA improve in the private setting compared to 1-SPSA. We leave the improvement in tighter analysis for private n-SPSA as future work and use 1-SPSA to conduct experiments.
@@
As discussed in Section 3Algorithm 1 can be incorporated in different noise mechanisms. We focus on the Gaussian noise mechanism and the Laplace mechanism in this work. The Gaussian noise mechanism has been widely studied in previous literature both for privacy analysis and empirical performance in DP-SGD (Abadi et al., 2016; Mironov, 2017; Dong et al., 2019). The Laplace mechanism, though less studied for privacy-preserving machine learning, can provide pure DP while the Gaussian mechanism can only provide approximate DP. We have provided the privacy analysis in Section A.
@@ Appendix D Hyperparameter Search
We have presented the hyperparameter search for batch size in Table 7. In this section, we first provide several initial results for hyperparameter search on clipping threshold and finally present the hyperparameter tables. We also provide an initial study to systematically evaluate the interplay between batch size and training iterations for DP-ZO.
@@
Li et al. (2022b); De et al. (2022) recommend small clipping $C$ threshold for DP-SGD training. For example, Li et al. (2022b) use $C=0.1$ for training language models. We therefore study the effect of different clipping threshold and present the results in Table 10. We find that while $C=1$ performs significantly worse, setting $C$ as $0.1$, $0.05$, $0.01$ are within the 2% performance gap. We therefore choose $C=0.05$.
@@
We present the hyperparameter for DP-ZO (Gaussian) in Table 11.
@@
learning rate [2e-5, 5e-5, 1e-4, 2e-4, 5e-4] for LoRA finetuning.
[2e-7, 5e-7, 1e-6, 2e-6, 5e-6] for Full fine-tuning.
LoRA rank 8
@@
In Fig. 5 and Fig. 5, we did an initial study to systematically evaluate the interplay between batch size and training iterations by varying batch size in [4,8,16,32,64] and steps in [10000, 2000]. Similar to main results, we run 5 independent runs for each setting and compute the average of 5 runs.
Figure 4: Ablationg ZO with Batch size and steps $T$. SQuAD dataset with 1000 training samples.
Figure 4: Ablationg ZO with Batch size and steps $T$. SQuAD dataset with 1000 training samples.
Figure 5: Ablationg DP-ZO with Batch size and steps $T$. SQuAD dataset with 1000 training samples. $(1,10^{-5})$-DP.
@@ Appendix E Ablation on Model Size
Section 4.2 shows that DP-ZO scales to larger models and provides the results of DP-ZO for model size varying from 1.3B to 66B parameters in Table 2. Here we provide the full results of DP-ZO finetuned with LoRA at $\varepsilon=1$, with model size ranging from 125M to 66B. We also include the $\varepsilon=0$ and $\varepsilon=\infty$ baseline as a reference in Table 13.
Table 13 shows the full trend of DP-ZO with model size scaling from 125M to 66B, that is DP-ZO scales to larger models.
For OPT-125m, the gap between private and non-private baseline is $4.96$. For OPT-66B, the non-private baseline is 85.73 and the gap between the private and non-private results is $1.86$.
@@
Model OPT-125M OPT-350M OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B OPT-66B
$\varepsilon=0$ $9.7$ $12.50$ $27.20$ $29.89$ $36.48$ $46.23$ $46.53$ $48.13$
$\varepsilon=1$ ${43.74_{1.49}}$ $50.69_{0.92}$ $74.41_{0.81}$ $79.30_{0.93}$ $79.53_{0.65}$ $81.35_{0.38}$ $82.03_{0.90}$ $83.87_{0.99}$
$\varepsilon=\infty$ $48.7$ $54.05$ $76.15$ $81.47$ $82.78$ $84.73$ $85.48$ $85.73$

### v2

@@ Appendix B Implementation Details
We follow Malladi et al. [2023] and provide the memory-efficient version of DP-ZO in Algorithm 2. Algorithm 2 enjoys the benefit that it does not incur additional GPU memory cost compared to inference.
Algorithm 2 Differentially Private-ZO (GPU memory efficient version. Adapted from Malladi et al. [2023])
@@
2:procedure DP-ZO(($\theta$, $\mathcal{D}$, $\epsilon$, $\sigma$, $T$, $s$, $\phi$, $C$, $\alpha$))
@@
17:         end for
18:     end for
19:end procedure
20:procedure PerturbParameters(($\theta,\phi,s$))
21:     Reset random number generator with seed $s$
22:     for $\theta_{i}\in\theta$ do
23:         $z\sim\mathcal{N}(0,1)$
24:         $\theta_{i}\leftarrow\theta_{i}+\phi z$
25:     end for
26:end procedure
@@ Appendix C Design Choices
$g_{i}$ $=L(\theta+\epsilon z_{i};B)-L(\theta-\epsilon z_{i};B)\text{(projected gradient for each }i)$
@@
Algorithm 1 estimates the gradients once. As outlined above, SPSA can be extended to n-SPSA gradient estimator and n-SPSA can improve the performance in the non-private setting [Malladi et al., 2023]. Here we discuss our design choice of why we choose $n=1$ in Algorithm 1.
  • Estimate the average. Previous work [Malladi et al., 2023] shows that averaged estimation helps the non-private setting. In a private setting, we have to privatize the gradient estimation. Here we discuss our initial design of the privatized n-SPSA gradient estimation. For the sampled batch, assuming we are adding the Gaussian noise $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ for 1-SPSA. Then for n-SPSA, to ensure we have the same privacy cost as 1-SPSA, we need to add $\mathcal{N}(0,n\cdot\mathrm{C}^{2}\sigma^{2})$ to each gradient estimation and finally average the $n$ gradients. Our privacy analysis follows the $n$-fold composition of Gaussian mechanism (Corollary 3.3 in Gaussian differential privacy [Dong et al., 2019]). Our initial experiment result shows that our current analysis for n-SPSA noise addition does not make n-SPSA improve in the private setting compared to 1-SPSA. We leave the improvement in tighter analysis for private n-SPSA as future work and use 1-SPSA to conduct experiments.
@@
As discussed in Section 3Algorithm 1 can be incorporated in different noise mechanisms. We focus on the Gaussian noise mechanism and the Laplace mechanism in this work. The Gaussian noise mechanism has been widely studied in previous literature both for privacy analysis and empirical performance in DP-SGD [Abadi et al., 2016, Mironov, 2017, Dong et al., 2019]. The Laplace mechanism, though less studied for privacy-preserving machine learning, can provide pure DP while the Gaussian mechanism can only provide approximate DP. We have provided the privacy analysis in Section A.
@@ Appendix D Hyperparameter Search
Our experiments are based on the open-source code [^2] https://github.com/princeton-nlp/MeZO. of Malladi et al. [2023]. We provide the prompts we use in Table Table 13. In this section, we first provide several initial results for hyperparameter search on clipping threshold and finally present the hyperparameter tables. We also provide an initial study to systematically evaluate the interplay between batch size and training iterations for DP-ZO.
Table 13: The prompts of the datasets we used for DP-ZO.
Dataset Type Prompt
SQuAD QA Title: <title>
Context: <context>
Question: <question>
Answer:
DROP QA Passage: <context>
Question: <question>
Answer:
SST-2 classification <text> It was terrible/great
@@
Li et al. [2022b], De et al. [2022] recommend small clipping $C$ threshold for DP-SGD training. For example, Li et al. [2022b] use $C=0.1$ for training language models. We therefore study the effect of different clipping threshold and present the results in Table 14. We find that while $C=1$ performs significantly worse, setting $C$ as $0.1$, $0.05$, $0.01$ are within the 2% performance gap. We therefore choose $C=0.05$.
@@
We present the hyperparameter for DP-ZO (Gaussian) in Table 15 and Table 16.
@@
$|\mathcal{D}|$ 1000
Steps $T$ $75000$
Clipping $C$ $0.05$
Batch size $16$
$\sigma$ $30.9$ for $\varepsilon=0.5$, $16.4$ for $\varepsilon=1$, $4.8$ for $\varepsilon=4$
learning rate [5e-6, 1e-5, 2e-5, 5e-5, 1e-4]
LoRA rank 8
Table 16: Hyperparameter search for DP-ZO with full parameter fine-tuning in Table 2.
@@
learning rate [2e-7, 5e-7, 1e-6, 2e-6, 5e-6]
@@
In Table 19 and Table 19, we did an initial study to systematically evaluate the interplay between batch size and training iterations by varying batch size in [16,32,64,128] and steps in [10000, 2000, 40000, 80000]. Similar to main results, we run 5 independent runs for each setting and compute the average of 5 runs. This ablation is by OPT-13B on SQuAD dataset with LoRA fine-tuning. Table 19 and Table 19 show that increasing steps $T$ improves the performance more than increasing the batch size. We also did ablation study on $T$ in [200, 400, 800, 1600] for DP-SGD (and did not observe significant improvements in DP-SGD) to ensure the fair comparison of DP-SGD and DP-ZO. Taking the computation limitation into consideration, we set $T=75000$ and BSZ=16 for main results in Table 1. We leave more investigation on the batch size and steps for DP-ZO, such as variance reduction method, as future work.
Table 18: $T=10000$, Varying batch size.
BSZ=16 BSZ=32 BSZ=64 BSZ=128
F1 81.35 81.63 81.47 81.72
Table 19: Batch size=16. Varying steps $T$.
$T$ $10000$ $20000$ $40000$ $80000$
F1 81.35 81.65 81.42 82.52
Computation Cost.
DP-ZO for OPT-13B models on SQuAD datasets takes around 4hrs for $10000$ steps. DP-SGD for OPT-13B models on SQuAD datasets takes around 4hrs for $200$ steps. When increasing $T$ or $B$ in DP-ZO, the training time scales proportionally to the scaling factor. Future work includes how to reduce the computation time of DP-ZO, e.g., by variance reduction method to improve the convergence rate.
@@ Appendix E Ablation on Model Size
Section 4.2 shows that DP-ZO scales to larger models and provides the results of DP-ZO for model size varying from 1.3B to 66B parameters in Table 2. Here we provide the full results of DP-ZO finetuned with LoRA at $\varepsilon=1$, with model size ranging from 1.3B to 66B. We also include the $\varepsilon=0$ and $\varepsilon=\infty$ baseline as a reference in Table 20.
Table 20 shows the full trend of DP-ZO with model size scaling from 1.3B to 66B, that is DP-ZO scales to larger models.
For OPT-1.3B, the gap between private and non-private baseline is $5.67$. For OPT-66B, the non-private baseline is $87.49$ and the gap between the private and non-private results is $3.37$.
@@
Model OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B OPT-66B
$\varepsilon=0$ $27.20$ $29.89$ $36.48$ $46.23$ $46.53$ $48.13$
$\varepsilon=1$ $75.29_{0.90}$ $80.34_{1.14}$ $81.34_{1.04}$ $82.28_{0.84}$ $82.48_{0.83}$ $84.12_{1.01}$
$\varepsilon=\infty$ $80.97$ $84.14$ $86.44$ $86.85$ $86.98$ $87.49$